OpenStudy (mackenzie2013):

What is the solution to the equation x/6= 4/12 x= 1 x= 2 x= 4 x= 6

jimthompson5910 (jim_thompson5910):

hint: cross multiply to get x/6 = 4/12 12x = 6*4 12x = 24 x = ???

OpenStudy (mackenzie2013):

2?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (mackenzie2013):

Thanks !

jimthompson5910 (jim_thompson5910):

check: x/6= 4/12 2/6= 4/12 ... plug in x = 2 1/3 = 4/12 1/3 = 1/3 ... this equation is true, so the answer is confirmed